Sheridan Inc. sells a product for $78 per unit. The variable cost is $44 per unit, while fixed costs are $298,248.

Determine (a) the break-even point in sales units and (b) the break-even point if the selling price were increased to $87 per unit.

a. Break-even point in sales units units
b. Break-even point if the selling price were increased to $87 per unit units   
Answer #1

a.Contribution margin=Sales-Variable cost

=(78-44)=$34 per unit

Breakeven=Fixed expenses/Contribution margin

=298,248/34

=8772 units

b.Contribution margin=Sales-Variable cost

=(87-44)=$43 per unit

Breakeven=Fixed expenses/Contribution margin

=298,248/43

=6936 units